# BiTrajDiff: Bidirectional Trajectory Generation with Diffusion Models for Offline Reinforcement Learning

### v2

@@ 1 Introduction
Offline Reinforcement Learning (Offline RL) [11] seeks to derive effective decision policies exclusively from pre-collected datasets, obviating the necessity for online exploration due to the inherent risks and costs associated with real-time interaction [25]. This paradigm has demonstrated notable efficacy across various real-world applications, including robotic control [38, 37, 29] and power grid management [3, 45]. A primary challenge in offline RL is addressing the out-of-distribution (OOD) problem, stemming from the inherent divergence between the static offline dataset distribution and the learned policy behavioral distribution during training. Consequently, the direct application of conventional online RL methods inevitably leads to extrapolation error [25], characterized by the overestimation of value functions for OOD state-action pairs. To tackle the OOD problem, offline RL research emphasizes the development of techniques that constrain learned policies to remain within the distribution of the static offline dataset. These approaches include imposing explicit constraint terms during policy optimization [38, 5] and penalizing the value function estimates for OOD actions [31, 46].
@@
Despite promising advancements in offline RL, current algorithms critically restrict agent actions, as evidenced by [25, 11, 38]. This limitation stems from the inherent mechanism constraining learned policy to the biased distribution of offline datasets, often leading to overfitting on limited behavioral patterns. While this principle effectively facilitates the replication offavorable behavioral patterns within the data, it inevitably curtails generalizability, hindering the exploration of optimal behaviors beyond dataset coverage. A common approach to mitigate the distribution bias is Data Augmentation (DA) method, which enriches the offline dataset with a large amount of synthesized high-fidelity data. For instance, studies like [30, 26, 27] leverage one generative diffusion model to approximate the distribution of single-direction forward-future trajectories for augmenting offline data. Notably, as data augmentation (DA) methods operate by modifying the training dataset, they are orthogonal to specific offline RL algorithms, enabling them to enhance any such algorithm without altering its core optimization mechanisms. However, a critical limitation of DA methods is their exclusive focus on trajectories originating from a predefined initial state. While these generated single-direction trajectories effectively explore behaviors emerging from a given state, they neglect the exploration of trajectories leading to that state. This oversight restricts the utility of the generated trajectory to locally explore the trajectory space immediately surrounding existing data, specifically for transitions originating from states already in the dataset. Consequently, while DA methods successfully generate trajectories for local state pairs already reachable within the offline dataset, they are ineffective in discovering global trajectories for previously unreachable state pairs. This naturally leads us to a neglected yet unresolved question:
@@
Our contributions are summarized as follows: (i). We are the first to explore and address the challenge of concurrently imagining both future and historical trajectories on a given state for data augmentation in offline RL. This paradigm enables the acquisition of diverse behavioral patterns. (ii). We propose BiTrajDiff, a novel framework that generates global trajectories between previously unreachable states in the original dataset, achieved by independently generating and then stitching together both forward-future and backward-history trajectories conditioned on common intermediate states. (iii). We conduct extensive experiments on the D4RL benchmark, applying BiTrajDiff across various offline RL algorithms. Our results demonstrate that BiTrajDiff achieves significantly superior performance compared to other advanced DA baselines, highlighting its effectiveness and robustness.
2 Related work
@@
Offline Reinforcement Learning encompasses four primary categories: policy constraint [44, 46], value regularization [23, 24], model-based [48, 2], and return-conditioned supervised learning [4, 34]. Policy constraint methods improve policies by maximizing expected returns while restricting agent actions to the offline dataset, employing techniques such as explicit behavior cloning regularization [38, 5] or implicit optimization towards derived in-sample optimal policies [32, 50]. Meanwhile, value regularization mitigates overestimation of OOD action values by promoting conservative value functions; recent work focuses on establishing precise lower bounds for Q-values to facilitate in-sample action selection [31, 33]. Model-based approaches construct task-oriented world models from offline data, including both one-step dynamics [48, 2] and multi-step sequence models [13, 14], which are subsequently used for offline reward penalization [49, 22] and online decision planning to ensure in-distribution actions [15, 16]. Return-conditioned supervised learning trains policies to generate trajectories conditioned on return values via supervised learning, with transformer [4, 39] and diffusion models [21, 1] serving as common policy backbones to enhance expressiveness and robustness. Nevertheless, a significant limitation of these approaches is their restriction of behavior to the inherent dataset distribution, which inevitably constrains the generalizability of the policies.
Data Augmentation (DA) Method for Offline RL leverages data augmentation to synthesize interaction data for subsequent RL training. To enhance data diversity and precision, Zhang et al. [51] train an ensemble of dynamics models and augment the RL buffer with generated transitions weighted by ensemble variance. Lu et al. [30] first utilize the notable generative capacity of diffusion to flexibly oversample generated one-step transitions for RL training. Subsequently, recent attempts in RL data augmentation have mainly focused on direct forward future trajectory generation. Several studies aim to achieve precise policy improvement by generating on-policy trajectories. For example, Jackson et al. [20] achieve on-policy trajectory generation by computing guidance derived from the policy model, whereas He et al. [17] employ interaction history embeddings to condition the generation process. Beyond on-policy generation, Wang et al. [43] employ diverse criteria, including dynamics error, temporal-difference error, and curiosity, to guide the augmented trajectory generation. Separately, Yang and Wang [47] investigate the generation of reverse trajectories terminated at conditioned states. Furthermore, Li et al. [26] explore stitching similar offline trajectories using intermediate states generated by masked diffusion. Nevertheless, a critical limitation of these DA methods is their inherent reliance on forward-only trajectory generation, which constrains them to local explorations around existing data. In contrast, our BiTrajDiff offers a distinct advantage by concurrently generating forward and backward trajectories starting from an intermediate state.
@@ 3 Preliminary
Reinforcement Learning (RL) paradigm [42] is formally defined as a Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,r,\gamma,\rho_{0}\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the environment dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1)$ is the discount factor, and $\rho_{0}$ denotes the initial state distribution. At each timestep $t$, an agent executes an action $a_{t}$ according to its policy $\pi(\cdot|s_{t})$ in the current state $s_{t}$, transitioning to the subsequent state $s_{t+1}$ with reward $r_{t}$ based on $P$. The objective of the agent is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted return: $\pi^{*}=\arg\max_{\pi}\mathcal{J}(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. In contrast to online RL, offline RL [25] constrains the agent from interaction with the environment, relying solely on a static trajectories dataset $\mathcal{D}=\{\tau^{i}\}_{i=1}^{N}$, where NN is the number of trajectories and each trajectory $\tau^{i}=\{(s^{i}_{t},a^{i}_{t},r^{i}_{t})\}_{t=0}^{T-1}$ comprises a sequence of state-action-reward tuples.
Diffusion Models [19, 41] are generative frameworks that learn data distributions through a forward diffusion process and a reverse denoising process. The forward process gradually corrupts data $\mathbf{x}_{0}$ via a predefined noise schedule $\{\beta_{k}\}_{k=1}^{K}$, where each transition is defined as $q(\mathbf{x}_{k}|\mathbf{x}_{k-1})=\mathcal{N}\left(\mathbf{x}_{k};\sqrt{1-\beta_{k}}\mathbf{x}_{k-1},\beta_{k}\mathbf{I}\right)$. As $K\to\infty$, the terminal distribution $q(\mathbf{x}_{K})$ converges to an isotropic Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$. The reverse process, parameterized by $\theta$, aims to iteratively reconstruct the original data through leanerable Gaussian transitions: $p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})=\mathcal{N}\left(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},k),\Sigma_{\theta}(\mathbf{x}_{k},k)\right).$ By maximizing the Empirical Lower Bound (ELBO) [40] on the log-likelihood of the sampled data, the diffusion model can be trained with a simplified surrogate loss [40]:
@@ 4 Methodology
This section introduces the BiTrajDiff framework for offline RL, which comprises two core components: bidirectional diffusion training and bidirectional trajectory generation. During the bidirectional diffusion training phase, two distinct diffusion models are employed to model the distributions of forward-future and backward-history state sequences within trajectory data. Conditioned on an intermediate state and cumulative reward signals, these models generate state sequences in corresponding temporal directions. Then, in the subsequent bidirectional trajectory generation phase, the generated forward and backward trajectories are reconciled through filling and filtering operations to produce novel global trajectories that connect previously unreachable states in the original dataset.
@@ 4.1 Bidirectional Diffusion Training
Building upon recent advances in diffusion-based sequence generation [8, 12], we propose a bidirectional trajectory generation framework comprising independent forward and backward trajectory diffusion models. These models are separately trained on offline datasets to learn the trajectory distribution of the behavior policy, facilitating bidirectional state trajectory generation from shared intermediate “anchor” states.
Formally, for each diffusion model, the state trajectory is generated simultaneously per time step $t$ over the planning horizon $H$: $\mathbf{x}_{k}(\tau)=\{s_{t},s_{t+1},\cdots,s_{t+H-1}\}_{k}$, where $k$ denotes the denoise timestep and $t$ represents the MDP timestep. In this way, the conditional generative problem via diffusion model can be formulated as:
$\min_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{\tau\sim\mathcal{D}}[\log p_{\theta}(\mathbf{x}_{0}(\tau)|\mathbf{y}(\tau))],$ (2)
where $\mathbf{x}_{0}(\tau)$ denotes the final generated subtrajectory of diffusion, $\mathbf{y}(\tau)$ represents the condition for generation, and $\theta$ represents the learnable parameters of the diffusion model. While various condition metrics are available, such as temporal difference error, Q-function, or cumulative return [43], our framework utilizes the cumulative return $R$ and the state $s$ as conditional input $\mathbf{y}(\tau)$. This approach, consistent with prior works [1, 6], enables the generation of trajectories with desired returns starting from the specific states. However, to enable bidirectional trajectory generation, the formations of the backward and forward diffusion condition inputs are quite distinct.
@@
Based on the two distinct conditions, we can model the conditional trajectory generation processes by the corresponding diffusion sampling composed of the forward noising process $q(\mathbf{x}_{k+1}(\tau)|\mathbf{x}_{k}(\tau))$ and the reverse denoising process $p_{\theta}(\mathbf{x}_{k-1}(\tau)|\mathbf{x}_{k}(\tau),\mathbf{y}(\tau))$ constructed with the diffusion denoise model $\epsilon_{\theta}$. Meanwhile, the trainable forward and backward diffusion models $\epsilon_{\theta}$ in our BiTrajDiff framework are parameterized with $\theta^{b}$ and $\theta^{f}$ respectively. Thus, given the condition $\mathbf{y}(\tau)$, the forward and backward diffusion models are able to generate the denoised trajectory $\mathbf{x}_{0}(\tau)$. This process begins with sampling random noise $\mathbf{x}_{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, followed by iterative denoising steps where $\mathbf{x}_{k-1}\sim p_{\theta}$, for $k=\{K,\cdots,1\}$.
To realize conditional diffusion model training and inference, we take the choice of Classifier-Free Guidance (CFG) technique [18], which uses only a single denoise model to handle both conditional and unconditional generation. Formally, the perturbed noise after CFG can be represented as:
@@
where the scalar $\omega\in[0,1]$ denotes the guidance weight of CFG. Setting $\omega$ to 0 degrades the generation to unconditional generation. Conversely, an increased value of $\omega$ promotes a stronger adherence of the generated samples to the provided condition $\mathbf{y}(\tau)$. To train such diffusion denoise models with CFG, the diffusion denoisers are trained with an objective that enables it to predict noise conditioned on $\mathbf{y}(\tau)$ and also unconditionally. This is typically achieved by randomly dropping the condition $\mathbf{y}(\tau)$ during training with probability pp. Specifically, we have the following loss:
@@
where $\text{Bern}(p)$ represents the Bernoulli distribution. In our implementation, we fix the state condition $s_{t}$ in $\mathbf{x}_{k}(\tau)$ during the adding-noise training as well as the denoising inference stage for simplicity and efficiency following [1, 21, 7, 47], while the diffusion denoise models $\epsilon_{\theta}$ only take the cumulative return $R(\tau)$ components of $\mathbf{y}(\tau)$ as condition inputs.

### v3

@@ 1 Introduction
Offline Reinforcement Learning (Offline RL) (fujimoto2019off) seeks to derive effective decision policies exclusively from pre-collected datasets, obviating the necessity for online exploration due to the inherent risks and costs associated with real-time interaction (levine2020offline). This paradigm has demonstrated notable efficacy across various real-world applications, including robotic control (qing2024a2po, qing2022XRLsurvey, liu2023curricular) and power grid management (chen2024powerformer, xu2024TPA). A primary challenge in offline RL is addressing the out-of-distribution (OOD) problem, stemming from the inherent divergence between the static offline dataset distribution and the learned policy behavioral distribution during training. Consequently, the direct application of conventional online RL methods inevitably leads to extrapolation error (levine2020offline), characterized by the overestimation of value functions for OOD state-action pairs. To tackle the OOD problem, offline RL research emphasizes the development of techniques that constrain learned policies to remain within the distribution of the static offline dataset. These approaches include imposing explicit constraint terms during policy optimization (qing2024a2po, chen2022lapo) and penalizing the value function estimates for OOD actions (lyu2022mildly, xu2023offline).
@@
Despite promising advancements in offline RL, current algorithms critically restrict agent actions, as evidenced by (levine2020offline, fujimoto2019off, qing2024a2po). This limitation stems from the inherent mechanism constraining learned policy to the biased distribution of offline datasets, often leading to overfitting on limited behavioral patterns. While this principle effectively facilitates the replication of favorable behavioral patterns within the data, it inevitably curtails generalizability, hindering the exploration of optimal behaviors beyond dataset coverage. A common approach to mitigate the distribution bias is Data Augmentation (DA) method, which enriches the offline dataset with a large amount of synthesized high-fidelity data. For instance, studies like (lu2023synthetic, li2024diffstitch, li2024augmenting) leverage one generative diffusion model to approximate the distribution of single-direction forward-future trajectories for augmenting offline data. Notably, as data augmentation (DA) methods operate by modifying the training dataset, they are orthogonal to specific offline RL algorithms, enabling them to enhance any such algorithm without altering its core optimization mechanisms. However, a critical limitation of DA methods is their exclusive focus on trajectories originating from a predefined initial state. While these generated single-direction trajectories effectively explore behaviors emerging from a given state, they neglect the exploration of trajectories leading to that state. This oversight restricts the utility of the generated trajectory to locally explore the trajectory space immediately surrounding existing data, specifically for transitions originating from states already in the dataset. Consequently, while DA methods successfully generate trajectories for local state pairs already reachable within the offline dataset, they are ineffective in discovering global trajectories for previously unreachable state pairs. This naturally leads us to a neglected yet unresolved question:
@@
Our contributions are summarized as follows: (1) We are the first to explore and address the challenge of concurrently imagining both future and historical trajectories on a given state for data augmentation in offline RL. This paradigm enables the acquisition of diverse behavioral patterns. (2) We propose BiTrajDiff, a novel framework that generates global trajectories between previously unreachable states in the original dataset, achieved by independently generating and then stitching together both forward-future and backward-history trajectories conditioned on common intermediate states. (3) We conduct extensive experiments on the D4RL benchmark, applying BiTrajDiff across various offline RL algorithms. Our results demonstrate that BiTrajDiff achieves significantly superior performance compared to other advanced DA baselines, highlighting its effectiveness and robustness.
2 Related Works
@@
Offline Reinforcement Learning encompasses four primary categories: policy constraint (wang2022diffusion, xu2023offline), value regularization (kostrikov2021offline, kumar2020conservative), model-based (yu2020mopo, chemingui2024offline), and return-conditioned supervised learning (chen2021decision, paster2022you). Policy constraint methods restrict policies to the offline dataset, employing techniques such as explicit behavior cloning regularization (qing2024a2po, chen2022lapo) or implicit optimization towards in-sample optimal policies (nair2020awac, yue2022boosting). Meanwhile, value regularization promotes conservative value functions to alleviate OOD overestimation via establishing precise lower bounds for Q-values to facilitate in-sample action selection (lyu2022mildly, park2024value). Model-based approaches construct task-oriented world models from offline data, including one-step dynamics (yu2020mopo, chemingui2024offline) and multi-step sequence models (hafner2020mastering, hafner2023mastering), which are subsequently used for offline reward penalization (yu2021combo, kidambi2020morel) and online decision planning to ensure in-distribution actions (hansen2022temporal, hansen2023td). Return-conditioned supervised learning trains trajectory generators conditioned on returns, with transformers (chen2021decision, schmied2024retrieval) and diffusion models (janner2022planning, ajay2022conditional) as common backbones. Despite their effectiveness, these approaches are fundamentally limited by reliance on the dataset distribution.
Data Augmentation for Offline RL synthesizes additional interactions to enrich the dataset for subsequent RL training. Early work employs dynamics ensembles to generate transitions with uncertainty weighting (zhang2023uncertainty) or diffusion models for flexible one-step oversampling (lu2023synthetic). More recent efforts shift toward trajectory-level generation. Some work focuses on on-policy rollouts, guided either by policy outputs (jackson2024policy) or interaction histories (he2023diffusion). Other approaches explore prioritized augmentation based on dynamics error or curiosity (wang2024prioritized), as well as trajectory recomposition through reverse generation (yang2025rtdiff) or diffusion-based stitching (li2024diffstitch). Despite these advances, existing DA methods remain limited by their forward-only generation paradigm, restricting exploration to local variations around the dataset. Our BiTrajDiff overcomes this limitation by jointly generating forward and backward trajectories from intermediate states.
@@ 3 Preliminary
Reinforcement Learning paradigm (sutton2018reinforcement) is formally defined as a Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,r,\gamma,\rho_{0}\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the environment dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1)$ is the discount factor, and $\rho_{0}$ denotes the initial state distribution. At each timestep $t$, an agent executes an action $a_{t}$ according to its policy $\pi(\cdot|s_{t})$ in the current state $s_{t}$, transitioning to the subsequent state $s_{t+1}$ with reward $r_{t}$ based on $P$. The objective of the agent is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted return: $\pi^{*}=\arg\max_{\pi}\mathcal{J}(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. In contrast to online RL, offline RL (levine2020offline) relys solely on a static trajectories dataset $\mathcal{D}=\{\tau^{i}\}_{i=1}^{N}$, where $N$ is the number of trajectories and each trajectory $\tau^{i}=\{(s^{i}_{t},a^{i}_{t},r^{i}_{t})\}_{t=0}^{T-1}$ comprises a sequence of state-action-reward tuples.
Diffusion Models (ho2020denoising, song2021maximum) are generative frameworks that learn data distributions through a forward diffusion process and a reverse denoising process. The forward process gradually corrupts data $\mathbf{x}_{0}$ via a predefined noise schedule $\{\beta_{k}\}_{k=1}^{K}$, where each transition is defined as $q(\mathbf{x}_{k}|\mathbf{x}_{k-1})=\mathcal{N}\left(\mathbf{x}_{k};\sqrt{1-\beta_{k}}\mathbf{x}_{k-1},\beta_{k}\mathbf{I}\right)$. As $K\to\infty$, the terminal distribution $q(\mathbf{x}_{K})$ converges to an isotropic Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$. The reverse process, parameterized by $\theta$, aims to iteratively reconstruct the original data through learnable Gaussian transitions: $p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})=\mathcal{N}\left(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},k),\Sigma_{\theta}(\mathbf{x}_{k},k)\right).$ By maximizing the Empirical Lower Bound (ELBO) (sohn2015learning) on the log-likelihood of the sampled data, the diffusion model can be trained with a simplified surrogate loss (sohn2015learning):
@@
Figure 1: An illustrative diagram of our Bidirectional Trajectory Diffusion (BiTrajDiff) method.
@@ 4 Methodology
This section introduces the BiTrajDiff framework for offline RL, which comprises two core components: bidirectional diffusion training and bidirectional trajectory generation. During the bidirectional diffusion training phase, two distinct diffusion models are employed to model the distributions of forward-future and backward-history state sequences within trajectory data. Conditioned on an intermediate state and cumulative reward signals, these models generate state sequences in corresponding temporal directions. Then, in the subsequent bidirectional trajectory generation phase, the generated forward and backward trajectories are reconciled through filling and filtering operations to produce novel global trajectories that connect previously unreachable states in the original dataset. The framework of our method is presented in Figure 1.
@@ 4.1 Bidirectional Diffusion Training
Building upon recent advances in diffusion-based sequence generation (fathi2025unifying, gong2022diffuseq), we propose a bidirectional trajectory generation framework comprising independent forward and backward trajectory diffusion models. These models are separately trained on offline datasets to learn the trajectory distribution of the behavior policy, facilitating bidirectional state trajectory generation from shared intermediate “anchor” states.
Formally, for each diffusion model, the state trajectory is generated simultaneously per time step $t$ over the planning horizon $H$: $\mathbf{x}_{k}(\tau)=\{s_{t},s_{t+1},\cdots,s_{t+H-1}\}_{k}$, where $k$ denotes the denoise timestep and $t$ represents the MDP timestep. In this way, the conditional generative problem via the diffusion model can be formulated as:
$\min_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{\tau\sim\mathcal{D}}[-\log p_{\theta}(\mathbf{x}_{0}(\tau)|\mathbf{y}(\tau))],$ (2)
where $\mathbf{x}_{0}(\tau)$ denotes the final generated subtrajectory of diffusion, $\mathbf{y}(\tau)$ represents the condition for generation, and $\theta$ represents the learnable parameters of the diffusion model. While various condition metrics are available, such as temporal difference error, Q-function (wang2024prioritized), our framework utilizes the cumulative return $R$ and the state $s$ as conditional input $\mathbf{y}(\tau)$. This approach, consistent with prior works (ajay2022conditional, ding2024diffusion), enables the generation of trajectories with desired returns starting from the specific states. However, to enable bidirectional trajectory generation, the formations of the backward and forward diffusion condition inputs are quite distinct.
@@
Based on the two distinct conditions, we can model the conditional trajectory generation processes by the diffusion sampling composed of the forward noising process $q(\mathbf{x}_{k+1}(\tau)|\mathbf{x}_{k}(\tau))$ and the reverse denoising process $p_{\theta}(\mathbf{x}_{k-1}(\tau)|\mathbf{x}_{k}(\tau),\mathbf{y}(\tau))$ constructed with the diffusion denoise model $\epsilon_{\theta}$. Meanwhile, the forward and backward diffusion models $\epsilon_{\theta}$ in our BiTrajDiff framework are parameterized with $\theta^{f}$ and $\theta^{b}$, respectively. Thus, given the condition $\mathbf{y}(\tau)$, they are able to generate the denoised trajectory $\mathbf{x}_{0}(\tau)$. This process begins with sampling random noise $\mathbf{x}_{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, followed by iterative denoising steps where $\mathbf{x}_{k-1}\sim p_{\theta}$, for $k=\{K,\cdots,1\}$.
To realize conditional diffusion model training and inference, we employ the Classifier-Free Guidance (CFG) technique (ho2022classifier), which uses a single denoise model to handle both conditional and unconditional generation. Formally, the perturbed noise after CFG can be formed as:
@@
where the scalar $\omega\in[0,1]$ denotes the guidance weight of CFG. Setting $\omega$ to 0 degrades the generation to unconditional generation. Conversely, an increased value of $\omega$ promotes a stronger adherence of the generated samples to the provided condition $\mathbf{y}(\tau)$. To train such diffusion denoise models with CFG, the diffusion denoisers are trained with an objective that enables them to predict noise conditioned on $\mathbf{y}(\tau)$ and also unconditionally. This is typically achieved by randomly dropping the condition $\mathbf{y}(\tau)$ during training with probability $p$. Specifically, we have the following loss function:
@@
where $\text{Bern}(p)$ represents the Bernoulli distribution. In our implementation, we fix the state condition $s_{t}$ in $\mathbf{x}_{k}(\tau)$ during the adding-noise training as well as the denoising inference stage for simplicity and efficiency following (ajay2022conditional, yang2025rtdiff), while the diffusion denoise models $\epsilon_{\theta}$ only take the cumulative return $R(\tau)$ components of $\mathbf{y}(\tau)$ as condition inputs.